OpenStudy (anonymous):

Andrew can clean a pool thrice as fast as Maggie. If both of them worked together for the first 3 hours, then Maggie stopped, it will require Andrew 4 hours more to finish the job. How many hours can Andrew clean the swimming pool alone?

OpenStudy (anonymous):

Can someone explain how to get the answer? Thanks!

OpenStudy (anonymous):

Hi, you can use the technique found here. Scroll down to the mechanic problem. http://www.purplemath.com/modules/workprob2.htm

OpenStudy (anonymous):

hi! i used the same technique used in the problem, however i'm getting a different answer. my equation is 7/3x+3/x=1. and i get x=16/3 where am i wrong?

OpenStudy (amistre64):

if your still there show me your steps and let me see if I can redress it

OpenStudy (anonymous):

7/3x+3/x=1 7+9=3x 16=3x 16/3=x

OpenStudy (amistre64):

the hiccup here is to account for maggie leaving after 3 hours and andrew having to finish up.

OpenStudy (amistre64):

work(hours) = 1 job; ex: (bobs.work)(5hours) = 1 job done bobs.work(1hour) = 1/5 job done A(3hrs) + M(3hrs) + A(4hrs) = 1 job done A(7hrs) + M(3hrs) = 1 job done

OpenStudy (amistre64):

Andrew is 3 times as fast as maggie; or is it 1/3 as fast? thrice means 3times i think

OpenStudy (amistre64):

M(3hrs) = A(1hr) A(7hrs) + A(1hr) 1 job done

OpenStudy (amistre64):

A(8hrs) = 1 job done

OpenStudy (amistre64):

3(1/8 + 3/8) +4(1/8) = 7 ?? 12/8 + 4/8 = ?? 16/8 = 2 hours tho; so I gotta wonder if im doing it right; or dbl chking it right :)

OpenStudy (anonymous):

hi! the answer key says the answer is 8.

OpenStudy (amistre64):

i guess i was correct then :) yay!!

OpenStudy (anonymous):

i finally got the solution right! i mixed up the rates and got confused with "Andrew can clean a pool thrice as fast as Maggie" 3/3x+7/x=1 3+21=3x 24=3x 8=x